# Benchmark on Drug Target Interaction Modeling from a Structure Perspective

### v1

@@ Appendix E Greedy Search on Hyperparameter
| GNN | GraphDTA-GCN [18] | 512 | 1000 | 0.0005 | - | 0.2 | Adam |
| GraphDTA-GAT [18] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| GraphDTA-GATGCN [18] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| GraphDTA-GIN [18] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| GraphCPI-GCN [35] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| GraphCPI-GAT [35] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| GraphCPI-GATGCN [35] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| GraphCPI-GIN [35] | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| MGraphDTA [59] | 512 | 292 | 0.0005 | - | 0.1 | Adam | |
| SAGDTA [60] | 512 | 2000 | 0.001 | - | 0.1 | Adam | |
| EmbedDTI [61] | 512 | 1500 | 0.0005 | - | 0.2 | Adam | |
| DeepGLSTM [29] | 128 | 1000 | 0.0005 | - | 0.2 | Adam | |
| CPI [31] | 1 | 100 | {0.001, 0.5, 10} | 1e-6 | 0 | Adam | |
| BACPI [57] | 16 | 20 | {0.0005, 0.5, 10} | - | 0.1 | Adam | |
| DeepNC-HGC [30] | 256 | 1000 | 0.0005 | - | 0.2 | Adam | |
| DeepNC-GEN [30] | 256 | 1000 | 0.0005 | - | 0.2 | Adam | |
| DrugBAN [62] | 64 | 100 | 0.0005 | - | 0 | Adam | |
| GANDTI [58] | 1 | 30/15 | 0.001 | 1e-6 | 0.5 | Adam | |
| PGraphDTA-CNN [24] | 512 | 1500 | 0.0005 | - | 0.2 | Adam | |
| BridgeDPI [63] | 512 | 100 | 0.001 | - | 0.5 | Adam | |
| ColdDTA [64] | 128 | 292 | 0.003 | - | 0 | Adam | |
| SubMDTA [36] | 512 | 1200 | 0.0005 | - | 0.2 | Adam | |
| IMAEN [65] | 128 | 1000 | 0.0005 | - | 0.2 | Adam | |
| Transformer | CSDTI [32] | 256 | 292 | 0.0005 | - | {0.5,0.2} | Adam |
| AMMVF [38] | 32 | 40 | {0.001, 0.5, 5} | 1e-4 | 0.1 | Adam | |
| TDGraphDTA [25] | 1024 | 3000 | 0.0005 | - | 0.1 | Adam | |
| IIFDTI [66] | 64 | 100 | 0.001 | 1e-6 | 0.2 | AdamW | |
| ICAN [42] | 128 | 50 | 0.001 | - | 0.1 | Adam | |
| MolTrans [20] | 16 | 13 | 0.001 | - | 0.1 | Adam | |
| TransformerCPI [21] | 8 | 40 | 0.0001 | 0.0001 | 0.2 | Radam | |
| MRBDTA [55] | 256 | 300 | 0.001 | - | 0.1 | Adam | |
| FOTFCPI [41] | 64 | 100 | 0.0001 | - | 0.1 | Adam |
| Models | Hyper parameter | Regression | Classification | |||||||||||||
| Batch size | Learning rate | Dropout | MSE | MAE | R2 | PCC | CI | Spearman | ROC-AUC | PR-AUC | Range-AUC | Acc. | Precision | Recall | F1 | |
| GraphDTA | 512 | 0.0005 | 0.2 | 0.2838 | 0.3053 | 0.5773 | 0.7619 | 0.8508 | 0.6226 | 0.9117 | 0.8807 | 0.4712 | 0.9117 | 0.9157 | 0.9125 | 0.9141 |
| 256 | 0.0005 | 0.2 | 0.2781 | 0.2924 | 0.5859 | 0.7677 | 0.8523 | 0.6250 | 0.9132 | 0.8805 | 0.4584 | 0.9135 | 0.9115 | 0.9218 | 0.9165 | |
| 128 | 0.0005 | 0.2 | 0.2634 | 0.2552 | 0.6077 | 0.7840 | 0.8700 | 0.6828 | 0.9134 | 0.8822 | 0.4716 | 0.9135 | 0.9157 | 0.9166 | 0.9160 | |
| 64 | 0.0005 | 0.2 | 0.2591 | 0.2506 | 0.6141 | 0.7885 | 0.8634 | 0.6862 | 0.9121 | 0.8793 | 0.4608 | 0.9123 | 0.9112 | 0.9195 | 0.9153 | |
| 512 | 0.0001 | 0.2 | 0.3209 | 0.3148 | 0.5221 | 0.7299 | 0.8369 | 0.6008 | 0.9165 | 0.8851 | 0.4775 | 0.9168 | 0.9159 | 0.9232 | 0.9195 | |
| 512 | 0.001 | 0.2 | 0.2909 | 0.3076 | 0.5668 | 0.7812 | 0.8425 | 0.6073 | 0.9133 | 0.8827 | 0.4765 | 0.9134 | 0.9172 | 0.9145 | 0.9157 | |
| 512 | 0.0005 | 0 | 0.2827 | 0.2923 | 0.5791 | 0.7632 | 0.8527 | 0.6267 | 0.9177 | 0.8863 | 0.4764 | 0.9180 | 0.9163 | 0.9252 | 0.9207 | |
| 512 | 0.0005 | 0.1 | 0.2771 | 0.2947 | 0.5873 | 0.7695 | 0.8521 | 0.6241 | 0.9170 | 0.8873 | 0.4830 | 0.9171 | 0.9204 | 0.9189 | 0.9194 | |
| 512 | 0.0005 | 0.5 | 0.2951 | 0.3165 | 0.5605 | 0.7494 | 0.8478 | 0.6176 | 0.9128 | 0.8819 | 0.4637 | 0.9129 | 0.9161 | 0.9151 | 0.9155 | |
| GraphCPI | 512 | 0.0005 | 0.2 | 0.3173 | 0.3309 | 0.5275 | 0.7325 | 0.8267 | 0.5825 | 0.9049 | 0.8687 | 0.4275 | 0.9053 | 0.9005 | 0.9177 | 0.9089 |
| 256 | 0.0005 | 0.2 | 0.2781 | 0.2924 | 0.5859 | 0.7677 | 0.8523 | 0.6250 | 0.9058 | 0.8712 | 0.4394 | 0.9061 | 0.9047 | 0.9140 | 0.9093 | |
| 128 | 0.0005 | 0.2 | 0.3064 | 0.2988 | 0.5436 | 0.7402 | 0.8467 | 0.6213 | 0.9085 | 0.8762 | 0.4602 | 0.9086 | 0.9115 | 0.9111 | 0.9112 | |
| 64 | 0.0005 | 0.2 | 0.3484 | 0.3569 | 0.4812 | 0.6944 | 0.8249 | 0.5850 | 0.9123 | 0.8774 | 0.4453 | 0.9128 | 0.9060 | 0.9270 | 0.9163 | |
| 512 | 0.0001 | 0.2 | 0.3461 | 0.3494 | 0.4846 | 0.7075 | 0.8317 | 0.5927 | 0.9063 | 0.8691 | 0.4217 | 0.9068 | 0.8985 | 0.9235 | 0.9107 | |
| 512 | 0.001 | 0.2 | 0.2870 | 0.2733 | 0.5726 | 0.7631 | 0.8598 | 0.6416 | 0.9064 | 0.8717 | 0.4438 | 0.9067 | 0.9048 | 0.9151 | 0.9099 | |
| 512 | 0.0005 | 0 | 0.3431 | 0.3527 | 0.4891 | 0.7096 | 0.8297 | 0.5890 | 0.9060 | 0.8709 | 0.4386 | 0.9064 | 0.9036 | 0.9160 | 0.9097 | |
| 512 | 0.0005 | 0.1 | 0.3291 | 0.3388 | 0.5100 | 0.7265 | 0.8294 | 0.5885 | 0.9060 | 0.8706 | 0.4385 | 0.9064 | 0.9029 | 0.9169 | 0.9098 | |
| 512 | 0.0005 | 0.5 | 0.3329 | 0.3264 | 0.5043 | 0.7304 | 0.8361 | 0.6211 | 0.9064 | 0.8702 | 0.4286 | 0.9068 | 0.9010 | 0.9203 | 0.9105 | |
| MRBDTA | 128 | 0.001 | 0.1 | 0.2397 | 0.2648 | 0.6430 | 0.8025 | 0.8716 | 0.6546 | 0.9340 | 0.9119 | 0.5240 | 0.9338 | 0.9427 | 0.9281 | 0.9353 |
| 64 | 0.001 | 0.1 | 0.2504 | 0.2679 | 0.6272 | 0.7935 | 0.8674 | 0.6483 | 0.9348 | 0.9113 | 0.5219 | 0.9348 | 0.9396 | 0.9336 | 0.9365 | |
| 32 | 0.001 | 0.1 | 0.2567 | 0.2721 | 0.6177 | 0.7878 | 0.8622 | 0.6400 | 0.9328 | 0.9127 | 0.5331 | 0.9324 | 0.9474 | 0.9200 | 0.9334 | |
| 128 | 0.0005 | 0.1 | 0.2350 | 0.2565 | 0.6499 | 0.8069 | 0.8775 | 0.6639 | 0.9382 | 0.9164 | 0.5304 | 0.9381 | 0.9442 | 0.9353 | 0.9397 | |
| 128 | 0.0001 | 0.1 | 0.2335 | 0.2696 | 0.6523 | 0.8084 | 0.8767 | 0.6630 | 0.9372 | 0.9171 | 0.5350 | 0.9370 | 0.9481 | 0.9284 | 0.9381 | |
| 128 | 0.001 | 0 | 0.2425 | 0.2663 | 0.6388 | 0.7997 | 0.8739 | 0.6587 | 0.9279 | 0.9010 | 0.4997 | 0.9280 | 0.9294 | 0.9316 | 0.9303 | |
| 128 | 0.001 | 0.2 | 0.2408 | 0.2620 | 0.6414 | 0.8027 | 0.8679 | 0.6478 | 0.9438 | 0.9218 | 0.5363 | 0.9438 | 0.9450 | 0.9460 | 0.9454 | |
| 128 | 0.001 | 0.5 | 0.2584 | 0.2666 | 0.6153 | 0.7894 | 0.8727 | 0.6573 | 0.9328 | 0.9123 | 0.5295 | 0.9325 | 0.9462 | 0.9215 | 0.9336 | |
| TransformerCPI | 128 | 0.001 | 0.1 | 0.3747 | 0.4111 | 0.4420 | 0.6787 | 0.8164 | 0.5668 | 0.9294 | 0.9024 | 0.5047 | 0.9296 | 0.9301 | 0.9339 | 0.9317 |
| 64 | 0.001 | 0.1 | 0.5916 | 0.5422 | 0.1191 | 0.3531 | 0.6644 | 0.3014 | 0.9318 | 0.9061 | 0.5100 | 0.9319 | 0.9337 | 0.9342 | 0.9339 | |
| 32 | 0.001 | 0.1 | 0.6219 | 0.5197 | 0.0740 | 0.2993 | 0.6360 | 0.2506 | 0.9269 | 0.8959 | 0.4860 | 0.9273 | 0.9198 | 0.9411 | 0.9302 | |
| 128 | 0.0005 | 0.1 | 0.2869 | 0.3389 | 0.5728 | 0.7750 | 0.8326 | 0.5910 | 0.9270 | 0.9008 | 0.5030 | 0.9270 | 0.9312 | 0.9270 | 0.9290 | |
| 128 | 0.0001 | 0.1 | 0.2877 | 0.3411 | 0.5716 | 0.7674 | 0.8348 | 0.5950 | 0.9208 | 0.8898 | 0.4664 | 0.9211 | 0.9181 | 0.9307 | 0.9239 | |
| 128 | 0.001 | 0 | 0.3874 | 0.3948 | 0.4232 | 0.6481 | 0.7985 | 0.5358 | 0.9297 | 0.9015 | 0.4999 | 0.9300 | 0.9272 | 0.9376 | 0.9324 | |
| 128 | 0.001 | 0.2 | 0.5181 | 0.5219 | 0.2284 | 0.5249 | 0.7476 | 0.4494 | 0.9236 | 0.8938 | 0.4780 | 0.9239 | 0.9217 | 0.9319 | 0.9265 | |
| 128 | 0.001 | 0.5 | 1.5318 | 1.1598 | -1.2809 | 0.2866 | 0.6441 | 0.2647 | 0.9218 | 0.8895 | 0.4746 | 0.9223 | 0.9153 | 0.9359 | 0.9253 | |
Appendix F Comparison of different featurization

### v2

@@ Appendix E Original Hyperparameter
| GNN | GraphDTA-GCN (nguyen2021graphdta) | 512 | 1000 | 0.0005 | - | 0.2 | Adam |
| GraphDTA-GAT (nguyen2021graphdta) | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| GraphDTA-GATGCN (nguyen2021graphdta) | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| GraphDTA-GIN (nguyen2021graphdta) | 512 | 1000 | 0.0005 | - | 0.2 | Adam | |
| GraphCPI-GCN (quan2019graphcpi) | 512 | 1000 | 0.0005 | - | 0.5 | Adam | |
| GraphCPI-GAT (quan2019graphcpi) | 512 | 1000 | 0.0005 | - | 0.6 | Adam | |
| GraphCPI-GATGCN (quan2019graphcpi) | 512 | 1000 | 0.0005 | - | - | Adam | |
| GraphCPI-GIN (quan2019graphcpi) | 512 | 1000 | 0.0005 | - | 0.6 | Adam | |
| MGraphDTA (yang2022MGraphDTA) | 512 | 3000 | 0.0005 | - | 0.1 | Adam | |
| SAGDTA (zhang2021sag) | 512 | 2000 | 0.001 | - | 0.1 | Adam | |
| EmbedDTI (jin2021embeddti) | 512 | 1500 | 0.0005 | - | 0.2 | Adam | |
| DeepGLSTM (mukherjee2022deepglstm) | 512/128 | 1000 | 0.0005 | - | 0.2 | Adam | |
| CPI (tsubaki2019compound) | 1 | 100 | 0.001, 0.5, 10 | 1e-6 | 0 | Adam | |
| BACPI (li2022bacpi) | 16 | 20 | 0.0005, 0.5, 10 | - | 0.1 | Adam | |
| DeepNC-HGC (tran2022deepnc) | 256 | 1000 | 0.0005 | - | 0.2 | Adam | |
| DeepNC-GEN (tran2022deepnc) | 256 | 1000 | 0.0005 | - | 0.2 | Adam | |
| DrugBAN (bai2023interpretable-drugban) | 64 | 100 | 0.00005 | - | 0 | Adam | |
| GANDTI (wang2021gandti) | 1 | 30/15 | 0.001 | 1e-6 | 0.5 | Adam | |
| PGraphDTA-CNN (bal2024pgraphdta) | 512 | 1500 | 0.0005 | - | 0.2 | Adam | |
| BridgeDPI (wu2022bridgedpi) | 512 | 100 | 0.001 | - | 0.5 | Adam | |
| ColdDTA (fang2023colddta) | 128 | 700/300 | 0.0003 | - | 0 | Adam | |
| SubMDTA (pan2023submdta) | 512 | 1200 | 0.0005 | - | 0.2 | Adam | |
| IMAEN (zhang2024imaen) | 128 | 1000 | 0.0005 | - | 0.2 | Adam | |
| Transformer | CSDTI (pan2023csdti) | 256 | 3000 | 0.0005 | - | 0.2 | Adam |
| AMMVF (wang2023ammvf) | 32 | 40 | 0.001, 0.5, 5 | 1e-4 | 0.1 | Adam | |
| TDGraphDTA (zhu2023tdgraph) | 1024 | 3000 | 0.0005 | - | 0.1 | Adam | |
| IIFDTI (cheng2022iifdti) | 64 | 200 | 0.001 | 1e-6 | 0.2 | AdamW | |
| ICAN (kurata2022ican) | 128 | 50 | 0.001 | - | 0.1 | Adam | |
| MolTrans (huang2021moltrans) | 64 | 30 | 0.00001 | - | 0.1 | Adam | |
| TransformerCPI (chen2020Transformercpi) | 8 | 40 | 0.0001, 0.5, 5 | 1e-4 | 0.2 | RAdam | |
| MRBDTA (zhang2022predicting) | 1024/256 | 600/300 | 0.001 | - | 0.1 | Adam | |
| FOTFCPI (yin2024fotf) | 64 | 100 | 0.0001 | - | 0.1 | Adam |
Appendix F Full experiment
To evaluate the model’s best performance, based on the hyperparameters given in its paper or codes, we found the optimized hyperparameters for each model. On top of the mean value, we also provide the standard deviation across five-fold. The complete result on the regression task is shown in Table 5, and the complete result on the classification task is shown in Table 6.
| Category | Models | DAVIS | KIBA | BindingDB | |||||||||||||||
| MSE | MAE | R2 | PCC | CI | Spearman | MSE () | MAE | R2 | PCC | CI | Spearman | MSE | MAE | R2 | PCC | CI | Spearman | ||
| GNN | GraphDTA-GCN | ||||||||||||||||||
| GraphDTA-GAT | |||||||||||||||||||
| GraphDTA-GATGCN | |||||||||||||||||||
| GraphDTA-GIN | |||||||||||||||||||
| GraphCPI-GCN | |||||||||||||||||||
| GraphCPI-GAT | |||||||||||||||||||
| GraphCPI-GATGCN | |||||||||||||||||||
| GraphCPI-GIN | |||||||||||||||||||
| MGraphDTA | |||||||||||||||||||
| SAGDTA | |||||||||||||||||||
| EmbedDTI | |||||||||||||||||||
| DeepGLSTM | |||||||||||||||||||
| CPI | |||||||||||||||||||
| BACPI | |||||||||||||||||||
| DeepNC-HGC | |||||||||||||||||||
| DeepNC-GEN | |||||||||||||||||||
| DrugBAN | |||||||||||||||||||
| GANDTI | |||||||||||||||||||
| BridgeDPI | |||||||||||||||||||
| ColdDTA | |||||||||||||||||||
| SubMDTA | |||||||||||||||||||
| IMAEN | |||||||||||||||||||
| Transformer | CSDTI | ||||||||||||||||||
| TDGraphDTA | |||||||||||||||||||
| AMMVF | |||||||||||||||||||
| IIFDTI | |||||||||||||||||||
| ICAN | - | - | |||||||||||||||||
| MolTrans | |||||||||||||||||||
| TransformerCPI | |||||||||||||||||||
| MRBDTA | |||||||||||||||||||
| FOTFCPI | |||||||||||||||||||
| Our combos | |||||||||||||||||||
| GNN | GraphDTA-GCN | |||||||||||||||||||||
| GraphDTA-GAT | ||||||||||||||||||||||
| GraphDTA-GATGCN | ||||||||||||||||||||||
| GraphDTA-GIN | ||||||||||||||||||||||
| GraphCPI-GCN | ||||||||||||||||||||||
| GraphCPI-GAT | ||||||||||||||||||||||
| GraphCPI-GATGCN | ||||||||||||||||||||||
| GraphCPI-GIN | ||||||||||||||||||||||
| MGraphDTA | ||||||||||||||||||||||
| SAGDTA | ||||||||||||||||||||||
| EmbedDTI | ||||||||||||||||||||||
| DeepGLSTM | ||||||||||||||||||||||
| CPI | ||||||||||||||||||||||
| BACPI | ||||||||||||||||||||||
| DeepNC-HGC | ||||||||||||||||||||||
| DeepNC-GEN | ||||||||||||||||||||||
| DrugBAN | ||||||||||||||||||||||
| GANDTI | ||||||||||||||||||||||
| BridgeDPI | ||||||||||||||||||||||
| ColdDTA | ||||||||||||||||||||||
| SubMDTA | ||||||||||||||||||||||
| IMAEN | ||||||||||||||||||||||
| Transformer | CSDTI | |||||||||||||||||||||
| TDGraphDTA | ||||||||||||||||||||||
| AMMVF | ||||||||||||||||||||||
| IIFDTI | ||||||||||||||||||||||
| ICAN | ||||||||||||||||||||||
| MolTrans | ||||||||||||||||||||||
| TransformerCPI | ||||||||||||||||||||||
| MRBDTA | ||||||||||||||||||||||
| FOTFCPI | ||||||||||||||||||||||
| Our combos | ||||||||||||||||||||||
Appendix G Comparison of different featurization